# ALMA: a mathematics-driven approach for determining tuning parameters in generalized LASSO problems, with applications to MRI

### v1

@@ Abstract
Magnetic Resonance Imaging (MRI) is a powerful technique employed for non-invasive in vivo visualization of internal structures. Sparsity is often deployed to accelerate the signal acquisition or overcome the presence of motion artifacts, improving the quality of image reconstruction. Image reconstruction algorithms use TV-regularized LASSO (Total Variation-regularized LASSO) to retrieve the missing information of undersampled signals, by cleaning the data of noise and while optimizing sparsity. A tuning parameter moderates the balance between these two aspects; its choice affecting the quality of the reconstructions. Currently, there is a lack of general deterministic techniques to choose these parameters, which are oftentimes manually selected and thus hinder the reliability of the reconstructions. Here, we present ALMA (Algorithm for Lagrange Multipliers Approximation), an iterative mathematics-inspired technique that computes tuning parameters for generalized LASSO problems during MRI reconstruction. We analyze quantitatively the performance of these parameters for imaging reconstructions via TV-LASSO in an MRI context on phantoms.
Although our study concentrates on TV-LASSO, the techniques developed here hold significant promise for a wide array of applications. ALMA is not only adaptable to more generalized LASSO problems but is also robust to accommodate other forms of regularization beyond total variation. Moreover, it extends effectively to handle non-Cartesian sampling trajectories, broadening its utility in complex data reconstruction scenarios. More generally, ALMA provides a powerful tool for numerically solving constrained optimization problems across various disciplines, offering a versatile and impactful solution for advanced computational challenges.
Index Terms:
Basis pursuit, compressed sensing, compressive sensing, convex analysis, convex optimization, denoising, inverse problems, LASSO, magnetic resonance, MRI, regression, sparsity, total variation
I Introduction
@@
If $\Phi$ is the $n\times n$ identity matrix, g-LASSO is called LASSO. In the case of MRI, where the sampled signal is the Fourier transform of the proton density weighted by the effects of T1 and T2 time-relaxations, cf. [1], $n$ is the number of voxels, $x$ is an image, $A=UFC$ encodes the undersampling operator $U$, the discrete Fourier transform (DFT) $F$ and $C$ is a complex matrix which encompasses the channel extension and the sensitivity mapping, $b$ is the noisy undersampled measurement and $\Phi$ can be the DFT, the discrete wavelet transform (DWT), the discrete cosine transform (DCT), or other sparsity promoting transforms, according to the situation, cf. [2]. The addendum $\|Ax-b\|_{2}^{2}$ measures the fidelity of the reconstruction, quantifying the distance between the measured data $b$ and the model $Ax$, whereas $\|\Phi x\|_{1}$ measures the sparsity of $\Phi x$. The parameter $\lambda>0$ acts as a trade-off, balancing the two contributions, and it is called tuning parameter. For small values of $\lambda$, the addendum $\lambda\|\Phi x\|_{1}$ becomes negligible and the image reconstructed via g-LASSO will be noisy and biased by the artifacts resulting from the missing information. On the contrary, large values of $\lambda$ force $\|\Phi x\|_{1}$ to be small, resulting in overly smoothed images with poor resolution (see fig. 1).
Figure 1: Impact of tuning parameters. Small values of $\lambda$ produce pronounced artifacts. Large values of $\lambda$ produce oversmoothing.
@@
The choice of this tuning parameter highly affects the quality of the reconstructions. However, there remains no generally effective procedure to provide well-performing tuning parameters. Consequently, the tuning parameter is chosen manually or heuristically, without following standard protocols. An automatic procedure to detect optimal tuning parameters is therefore needed to heavily reduce the post-processing of MR acquired signals, to increase the reproducibility of MR reconstructions, and the adequacy of reconstructions, according to the specific clinical or research question at hand.
@@
The mathematical motivation for considering g-LASSO problems relates to the constrained LASSO problem (or quadratically constrained basis pursuit):
$\text{minimize}\qquad\|x\|_{1}\ \text{subject to}\ x\in\mathbb{R}^{n},\quad\|Ax-b\|_{2}\leq\eta,$ (2)
@@
whose solution, if unique, is known to be $m$-sparse, where $m$ is the rank of the measurement matrix $A\in\mathbb{R}^{m\times n}$, cf. [3]. Under these assumptions, there exists a parameter $\lambda^{\prime}\geq 0$, related to the Lagrange multipliers, such that the unconstrained LASSO has an $m$-sparse solution. For simplicity, we refer to $\lambda^{\prime}$ as Lagrange multiplier. In broad terms, solving LASSO with tuning parameter $\lambda^{\prime}$ provides sparse solutions. The question thus arises whether Lagrange multipliers play an analogous role for generalized LASSO problems, where $\Phi$ is a general sparsity promoting transform. To answer this question, we build an iterative algorithm, that we call ALMA (Algorithm for Lagrange Multipliers Approximation), to provide an approximate Lagrange multipliers (ALM) and reconstruct a MR image from undersampled data, artificially corrupted by Gaussian noise, according to the MRI sampling procedure, by solving g-LASSO with such obtained parameter. The quality of the reconstructions, along with the optimality of the tuning parameters, is measured by means of three metrics, quantifying the main aspects of an effective MRI reconstruction: the impact of noise and the impact of artifacts, see Section III-C below. The purpose of this study is three-fold:
@@
  • To evaluate the quality of reconstructions by means of image quality metrics.
@@
II Theoretical model
II-A Mathematical rationale
The LASSO problem is fundamental in the recovery of sparse vectors. Its origins trace back to the seminal contributions of F. Santosa and W. W. Symes [4], S. Chen and D. Donoho [5, 6], and R. Tibshirani [7]. Notably, within the domain of MRI, M. Lustig and colleagues [2, 8, 9, 10, 11] pioneered the applications of both LASSO and g-LASSO. The purpose of this section is to motivate our conjecture that Lagrange multipliers could be used as effective tuning parameters for imaging reconstruction with g-LASSO. For the sake of completeness, we shall clarify the notion of sparsity comprehensively.
@@ II-A Mathematical rationale
$\text{minimize}\qquad\|x\|_{1}\ \text{subject to}\ x\in\mathbb{R}^{n},\ \|Ax-b\|_{2}\leq\eta$ (3)
@@
$A\in\mathbb{R}^{m\times n}$ is fundamental in Theorem 1, which fails if $A\in\mathbb{C}^{m\times n}$. For this reason, when working on $\mathbb{C}^{d}$, we consider its structure as the real vector space $\mathbb{R}^{2d}$.
@@
$\text{minimize}\qquad\frac{1}{2}\|Ax-b\|_{2}^{2}+\frac{\lambda}{2}\|\Phi x\|_{1},$ (5)
the rescaling by the factor $1/2$ is needed for computational purposes, but it is irrelevant to the analysis of (5). Also, observe that (5) with $\lambda\neq 0$ is equivalent to:
@@
II-B Construction of Lagrange multipliers
The construction of a Lagrange multiplier is detailed in the proof of [12, Theorem 4.8], which uses Hahn-Banach theorem to find a hyperplane that separates two convex sets, which for (6) read as: the epigraph $\mathcal{A}=\Big{\{}(u,t)\in\mathbb{R}^{2}:u\geq\frac{1}{2}\|Ax-b\|_{2}^{2}-\frac{\eta^{2}}{2},\ t\geq\frac{1}{2}\|\Phi x\|_{1}\ for\ some\ x\in\mathbb{R}^{n}\Big{\}}$ and the lower half-line =$\mathcal{B}=\Big{\{}(0,t)\in\mathbb{R}^{2}:t<p^{\ast}\Big{\}}$, where $p^{\ast}=\min\{\|\Phi x\|_{1}:\|Ax-b\|_{2}\leq\eta\}$ (see fig. 2).
@@ II-C Main challenges
While acknowledging the necessity of numerical analysis, we may resort to an escamotage to overcome the two remaining challenges. However, this entails abandoning the pursuit of exact Lagrange multipliers in favor of approximations. The constraint bound $\eta$ depends intrinsically on the solution $x^{\#}$ of g-LASSO, which in general differs from the ground truth $f$ which, in the case of our experimental setting, consists of the Shepp-Logan phantom (see fig. 5 (A)).
@@
II-D The MRI model
For the sake of concreteness, we simulate the reconstruction of a MR signal. Let us spend a few words about how g-LASSO is used, and why the correct choice of $\lambda$ is fundamental, in the context of MRI. Roughly speaking, the (inverse) spatial Fourier transform of the MR signal is the anatomical image of a tissue, which is supported in a cube $[-L_{x}/2,L_{x}/2]\times[-L_{y}/2,L_{y}/2]\times[-L_{z}/2,L_{z}/2]$. By Shannon’s theorem full Cartesian (uniform) sampling consists of sampling $n_{x}$ points in the $k_{x}$ direction, $n_{y}$ points along the $k_{y}$ direction and $n_{z}$ points along the $k_{z}$ direction of the k-space, where:
@@ II-D The MRI model
we mention [1] as reference therein.
For various reasons, sampling the MRI signal at its Nyquist frequency poses several challenges, necessitating techniques that can accurately reconstruct the MRI signal from samples taken below the Nyquist frequency. First, in our 2D experiments the dimensionality of the problem is of the order of $n=n_{x}n_{y}=384^{2}\sim 10^{5}$. For a 3D image, it increases to $n=n_{x}n_{y}n_{z}\sim 10^{8}$, making the processing of the MRI signal extremely time-consuming and computationally expensive. Secondly, MRI requires patients to remain still throughout the entire acquisition process, making it challenging to image moving organs and to perform scans on patients with conditions such as movement disorders. Furthermore, MRI is extremely expensive, and reducing the amount of information to be acquired can lead to significant cost savings.
In light of these reasons, CS offers a valuable solution for reducing both acquisition time and computational costs. However, there are situations where CS transcends being merely advantageous: it becomes imperative. Take, for example, 3D-CINE MRI, where sampling a full-Cartesian grid would be excessively time-consuming to the extent that achieving full sampling becomes practically infeasible. The application of CS to MRI is called CS-MRI. Since the MR signal is known to be sparse with respect to the DFT, DWT, and other sparsity promoting transforms, cf. [2], g-LASSO is used for the purpose. In this work, we use the Shepp-Logan phantom and corrupt the data with artificial noise. The Shepp-Logan phantom is piecewise constant and, therefore, its gradient is sparse. For this reason, we use the discrete gradient as sparsity promoting transform:
@@ III-A The simulated MR signal
    Channel extension and acquisition across coils: $f$ is replicated $nCh=8$ times and each replica $f_{j}\in\mathbb{R}^{384\times 384}$, $j=1,…,nCh$, is pointwise multiplied by a simulated coil sensitivity matrix $C_{j}\in\mathbb{C}^{384\times 384}$:
@@
    $k=1,…,nCh$, $i,j=1,…384$ (fig. 3).
    Figure 3: Simulated channel extension and acquisition across coils.
@@
    Fourier transform: the spatial discrete Fourier transform of each $f_{j}$ is computed (see fig. 9B).
@@
    Undersampling: full Cartesian sampling consists of sampling $nLines=384$ lines evenly spaced, whereas undersampling entails sampling only a certain fraction of these 384 lines, which we denote by ${UR}_{\%}$. In our study, we tested undersampling rates of $10\%$, $15\%$ and $20\%$, that is ${UR}_{\%}\in\{10/100,15/100,20/100\}$. For a fixed ${UR}_{\%}$, the sampling trajectory comprises $n({UR}_{\%})=\lceil nLines\cdot{UP}_{\%}\rceil$ lines. The $30\%$ of the $n({UR}_{\%})$ lines are used to sample the center of the k-space at the Nyquist frequency. Specifically, $\lceil n({UR}_{\%})\cdot 30/100\rceil$ lines sample the center of the k-space, while the remaining lines sample the periphery of the k-space following a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with $\mu=nLines/2+1$ and $\sigma^{2}=nLines\cdot{UR}_{\%}$ (see fig. 9C). The Fourier transform of each coil-image is sampled according to the Cartesian undersampled trajectory established before. Consequently, the resulting simulated MR data consists of a tensor $y\in Y=\mathbb{C}^{(nLines\cdot n({UR}_{\%}))\times nCh}$ (see fig. 9D).
@@
    Noise corruption: we tested ALMA under three distinct noise levels. Specifically, the simulated MR data is given by: $b=y+\varepsilon$, where $\varepsilon\in Y$ and $\Re(\varepsilon_{i,j}),\Im(\varepsilon_{i,j})\sim\mathcal{N}(0,\sigma^{2})$ ($i=1,…,nLines\cdot n({UR}_{\%}$) and $j=1,…,nCh$). We refer to $\sigma^{2}$ as to noise level, which is computed as $\sigma^{2}=\|y\|\cdot{NL}_{\%}$, where in our experiments ${NL}_{\%}\in\{3/100,5/100,7/100\}$. For instance, the terminology $3\%$ noise means that we are considering ${NL}_{\%}=3/100$.
III-B ALMA

### v2

@@ Abstract
Summary: ALMA (Algorithm for Lagrange Multipliers Approximation) is an iterative method for selecting tuning parameters in generalized LASSO problems for MRI reconstruction. Standardized approaches are often manual or heuristic; ALMA addresses this by approximating adaptatively Lagrange multipliers during reconstruction. On simulated MRI data, ALMA achieved mSSIM $\geq$ 0.99 and pSNR $\geq$ 40 dB, showing robust and near-optimal performance.
Purpose: To introduce ALMA, a method for adaptatively computing tuning parameters in generalized LASSO problems during MRI reconstruction.
Methods: We simulated MRI data using the Shepp-Logan phantom, testing ALMA across 450 reconstructions with varying undersampling (10–20%) and noise levels (3–7%). Image quality was assessed using mSSIM, pSNR, and CJV.
Results: ALMA achieved average mSSIM of $0.9951\pm 0.0041$, pSNR of $42.24\pm 3.42$ dB, and CJV of $0.0367\pm 0.0125$, converging in about $7.2\pm 3$ iterations.
Conclusion:ALMA offers a reliable, automated alternative to manual tuning, producing high-quality reconstructions and showing promise for in vivo MRI applications.
Keywords: Basis pursuit, compressed sensing, convex analysis, denoising, inverse problems, LASSO, MRI, sparsity, Total Variation.
1 Introduction
@@
If $\Phi$ is the $n\times n$ identity matrix, g-LASSO is called LASSO. In the case of MRI, where the sampled signal is the Fourier transform of the proton density weighted by the effects of T1 and T2 time-relaxations, cf. [1], $n$ is the number of voxels, $x$ is an image, $A=UFC$ encodes the undersampling operator $U$, the discrete Fourier transform (DFT) $F$ and $C$ is a complex matrix which encompasses the channel extension and the sensitivity mapping, $b$ is the noisy undersampled measurement and $\Phi$ can be the DFT, the discrete wavelet transform (DWT), the discrete cosine transform (DCT), or other sparsity promoting transforms, according to the situation, cf. [2]. The addendum $\|Ax-b\|_{2}^{2}$ measures the fidelity of the reconstruction, quantifying the distance between the measured data $b$ and the model $Ax$, whereas $\|\Phi x\|_{1}$ measures the sparsity of $\Phi x$. The parameter $\lambda>0$ acts as a trade-off, balancing the two contributions, and it is called tuning parameter. For small values of $\lambda$, the addendum $\lambda\|\Phi x\|_{1}$ becomes negligible and the image reconstructed via g-LASSO will be noisy and biased by the artifacts resulting from the missing information. On the contrary, large values of $\lambda$ force $\|\Phi x\|_{1}$ to be small, resulting in overly smoothed images with poor resolution (see fig.2B).
@@
The choice of this tuning parameter highly affects the quality of the reconstructions. However, there remains no generally effective procedure to provide well-performing tuning parameters. Consequently, the tuning parameter is chosen manually or heuristically, without following standard protocols. An automatic procedure to detect optimal tuning parameters is therefore needed to heavily reduce the post-processing of MR acquired signals, to increase the reproducibility of MR reconstructions, and the adequacy of reconstructions, according to the specific clinical or research question at hand. Of note, the tuning parameters can also be chosen by some statistical model, but those need to be trained on good quality fully sampled datasets. While such training dataset exist for images of low dimension (3D or less), high-dimensional images (4D or more) cannot be completely sampled in practice. Finding a good tuning parameter to reconstruct such images therefore still rely on methods that are free of statistical models, such as ALMA.
@@
The mathematical reason why considering g-LASSO problems relates to the constrained LASSO problem (or quadratically constrained basis pursuit):
$\text{minimize}\qquad\|x\|_{1}$ (2)
$\text{subject to}\ x\in\mathbb{R}^{n},\quad\|Ax-b\|_{2}\leq\eta,$
@@
whose solution, if unique, is known to be $m$-sparse, where $m$ is the rank of the measurement matrix $A\in\mathbb{R}^{m\times n}$, cf. [3]. Under these assumptions, there exists a parameter $\lambda^{\prime}\geq 0$, related to the Lagrange multipliers, such that the unconstrained LASSO has an $m$-sparse solution. For simplicity, we refer to $\lambda^{\prime}$ as Lagrange multiplier. In broad terms, solving LASSO with tuning parameter $\lambda^{\prime}$ provides sparse solutions. The question thus arises whether Lagrange multipliers play an analogous role for generalized LASSO problems, where $\Phi$ is a general sparsity promoting transform. To answer this question, we build an iterative algorithm, that we call ALMA (Algorithm for Lagrange Multipliers Approximation), to provide an approximate Lagrange multipliers (ALM) and reconstruct a MR image from undersampled data, artificially corrupted by Gaussian noise, according to the MRI sampling procedure, by solving g-LASSO with such obtained parameter. The quality of the reconstructions, along with the optimality of the tuning parameters, is measured by means of three metrics, quantifying the main aspects of an effective MRI reconstruction: the impact of noise and the impact of artifacts, see Section 2.2.3 below. The purpose of this study is three-fold:
@@
  • To assess the quality of reconstructions by means of image quality metrics.
@@
2 Methods
2.1 Theoretical model
2.1.1 Mathematical rationale
The LASSO problem is fundamental in the recovery of sparse vectors. It stemmed by the contributions of F. Santosa and W.W. Symes [4], S. Chen and D. Donoho [5, 6], and R. Tibshirani [7]. Notably, within the domain of MRI, M. Lustig and colleagues [2, 8, 9, 10, 11] pioneered the applications of LASSO and g-LASSO. The purpose of this section is to motivate our conjecture that Lagrange multipliers could be used as effective tuning parameters for imaging reconstruction with g-LASSO. For the sake of completeness, we shall clarify the notion of sparsity comprehensively.
@@ 2.1.1 Mathematical rationale
$\text{minimize}\qquad\|x\|_{1}$ (3)
$\text{subject to}\ x\in\mathbb{R}^{n},\ \|Ax-b\|_{2}\leq\eta$
@@
$A\in\mathbb{R}^{m\times n}$ is fundamental in Theorem 1, which fails if $A\in\mathbb{C}^{m\times n}$. For this reason, when working on $\mathbb{C}^{d}$, we consider its structure as the real vector space $\mathbb{R}^{2d}$ (see [12]).
@@
$\text{minimize}\qquad\frac{1}{2}\|Ax-b\|_{2}^{2}+\frac{\lambda}{2}\|\Phi x\|_{1}$ (5)
Note that the rescaling by the factor $1/2$ is due to a choice of implementation of the reconstruction, but it is irrelevant to the analysis of (5) from a mathematical point of view. Also, observe that (5) with $\lambda\neq 0$ is equivalent to:
@@
2.1.2 Construction of Lagrange multipliers
The construction of a Lagrange multiplier is detailed in the proof of [13, Theorem 4.8], which uses Hahn-Banach theorem to find a hyperplane that separates two convex sets (see fig.1), which for (6) read as the epigraph
[…]
$u\geq\frac{1}{2}\|Ax-b\|_{2}^{2}-\frac{\eta^{2}}{2},\ t\geq\frac{1}{2}\|\Phi x\|_{1},$
$\text{for some }x\in\mathbb{R}^{n}\Big{\}}$
and the lower half-line
=$\mathcal{B}=\Big{\{}(0,t)\in\mathbb{R}^{2}:t<p^{\ast}\Big{\}}$ (8)
where
$p^{\ast}=\min\{\|\Phi x\|_{1}:\|Ax-b\|_{2}\leq\eta\}$ (9)
@@ 2.1.3 Main challenges
While acknowledging the necessity of numerical analysis, we may resort to an escamotage to overcome the two remaining challenges. However, this entails abandoning the pursuit of exact Lagrange multipliers in favor of approximations. The constraint bound $\eta$ depends intrinsically on the solution $x^{\#}$ of g-LASSO, which in general differs from the ground truth $f$ which, in the case of our experimental setting, consists of the Shepp-Logan phantom (see fig.10C ).
@@
2.1.4 Comparison with the L-curve parameter
In this work, we compare the performance of ALMA with the L-curve method, a classical approach to determining and computing a suitable value for the regularization parameter $\lambda$ in (5). It was initially introduced by Lawson and Hanson in [15] and it consists of considering the curve:
$L=\{(\log\|Ax_{\lambda}-b\|_{2}^{2},\log\|\Phi x_{\lambda}\|_{1}),\quad\lambda>0\},$ (10)
where $x_{\lambda}$ is the solution of (5) with respect to the corresponding parameter $\lambda$. Hansen and O’Leary in [16, 17] proposed that the parameter corresponding to the corner of the graph (10) provides the best trade-off between $\|Ax-b\|_{2}^{2}$ and $\|\Phi x\|_{1}$.
2.1.5 The MRI model
For the sake of concreteness, we simulate the reconstruction of a MR signal. Let us spend a few words about how g-LASSO is used, and why the correct choice of $\lambda$ is fundamental, in the context of MRI. Roughly speaking, the (inverse) spatial Fourier transform of the MR signal is the anatomical image of a tissue, which is supported in a box $[-L_{x}/2,L_{x}/2]\times[-L_{y}/2,L_{y}/2]\times[-L_{z}/2,L_{z}/2]$. By Shannon’s theorem, full Cartesian (uniform) sampling consists of sampling $n_{x}$ points in the $k_{x}$ direction, $n_{y}$ points along the $k_{y}$ direction and $n_{z}$ points along the $k_{z}$ direction of the k-space, where:
@@ 2.1.5 The MRI model
We mention McRobbie et.al [1] as reference therein.
For various reasons, sampling the MRI signal at its Nyquist frequency poses several challenges. These challenges necessitate techniques that can accurately reconstruct the MRI signal from samples taken below the Nyquist frequency. First, in our 2D experiments the dimensionality of the problem is of the order of $n=n_{x}n_{y}=384^{2}\sim 10^{5}$. For a 3D image, it increases to $n=n_{x}n_{y}n_{z}\sim 10^{8}$, making the processing of the MRI signal extremely time-consuming and computationally expensive. Secondly, MRI requires patients to remain still throughout the entire acquisition process, making it challenging to image moving organs and to perform scans on patients with conditions such as movement disorders. Furthermore, MRI is extremely expensive, and reducing the amount of information to be acquired can lead to significant cost savings.
In light of these reasons, CS offers a valuable solution for reducing acquisition time. However, there are situations where CS transcends being merely advantageous: it becomes imperative. Take, for example, 3D-CINE MRI, where sampling a full-Cartesian grid would be excessively time-consuming to the extent that achieving full sampling becomes practically infeasible. The application of CS to MRI is called CS-MRI. Since the MR signal is known to be sparse with respect to the DFT, DWT, and other sparsity promoting transforms, cf. [2], g-LASSO is used for the purpose. In this work, we use the Shepp-Logan phantom and corrupt the data with artificial noise. The Shepp-Logan phantom is piecewise constant and, therefore, its gradient is sparse. For this reason, we use the discrete gradient as sparsity promoting transform:
@@ 2.2.1 The simulated MR signal
    Channel extension and acquisition across coils: $f$ is replicated $nCh=8$ times and each replica $f_{j}\in\mathbb{R}^{384\times 384}$, $j=1,\ldots,nCh$, is pointwise multiplied by a simulated coil sensitivity matrix $C_{j}\in\mathbb{C}^{384\times 384}$:
@@
    $k=1,\ldots,nCh$, $i,j=1,\ldots,384$ (fig. 2A).
    Figure 2: A: Simulated channel extension and pointwise multiplication by the coil-sensitivities. B: Reconstructions with different tuning parameters. For small values of $\lambda$, the reconstructed image exhibits pronounced artifacts. Large values of $\lambda$ produce oversmoothing.
@@
    Fourier transform: the spatial discrete Fourier transform of each $f_{j}$ is computed.
@@
    Undersampling: full Cartesian sampling consists of sampling $nLines=384$ lines evenly spaced, whereas undersampling entails sampling only a certain fraction of these 384 lines, which we denote by ${UR}_{\%}$. In our study, we tested undersampling rates of $10\%$, $15\%$ and $20\%$, that is ${UR}_{\%}\in\{10/100,15/100,20/100\}$. For a fixed ${UR}_{\%}$, the sampling trajectory comprises $n({UR}_{\%})=\lceil nLines\cdot{UP}_{\%}\rceil$ lines. The $30\%$ of the $n({UR}_{\%})$ lines are used to sample the center of the k-space at the Nyquist frequency. Specifically, $\lceil n({UR}_{\%})\cdot 30/100\rceil$ lines sample the center of the k-space, while the remaining lines sample the periphery of the k-space following a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with $\mu=nLines/2+1$ and $\sigma^{2}=nLines\cdot{UR}_{\%}$. The Fourier transform of each coil-image is sampled according to the Cartesian undersampled trajectory established before. Consequently, the resulting simulated MR data consists of a tensor $y\in Y:=\mathbb{C}^{(nLines^{2}\cdot n({UR}_{\%}))\times nCh}$.
@@
    Noise corruption: we tested ALMA under three distinct noise levels. Specifically, the simulated MR data is given by: $b=y+\varepsilon$, where $\varepsilon\in Y$ and $\Re(\varepsilon_{i,j}),\Im(\varepsilon_{i,j})\sim\mathcal{N}(0,\sigma^{2})$ ($i=1,\ldots,nLines\cdot n({UR}_{\%}$) and $j=1,\ldots,nCh$). We refer to $\sigma^{2}$ as to noise level, which is computed as $\sigma^{2}=\|y\|\cdot{NL}_{\%}$, where in our experiments ${NL}_{\%}\in\{3/100,5/100,7/100\}$. For instance, the terminology $3\%$ noise means that we are considering ${NL}_{\%}=3/100$.
2.2.2 ALMA